Billy goes to the store. He has $90. He wants to purchase a jacket for $45, a hat for $10, and the rest on jeans. Each pair of jeans costs $35. Write an inequality for the number of jeans he can purchase. If you are able to solve the problem, show your steps and solution as there are two steps to your solution. Use numbers and/or words in complete well-written sentences to explain your solution.
Answer by rothauserc(4718)

let x be the amount spent on jeans, then
1) 45 + 10 + x = 90
x = 90 - 55 = 35
:
2) let n be the number of jeans he will purchase, then
35n < or = x
Note that x = 35, therefore
35n < or = 35
n < or = 1
:
Billy has enough money left over from buying the jacket and hat to purchase 1 pair of jeans.